40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1254     Entry time: Wed Jan 28 12:42:51 2009
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: MOPA dying 
Yoichi, Jenne, Peter

As most of you know, the MOPA output power has been declining rapidly since Jan 21. (See the attachment 1)
There was also an increase in the NPRO power observed in LMON, which is an internal power monitor of the NPRO.
Similar trend can be seen in 126MON, which picks up some scattered light from the NPRO but there may be some contributions from the PA output.

The drop in the AMPMON, LMON and CURMON (NPRO current) from the middle of Jan 26 to the end of Jan27 was caused by me.
I tried to decrease the NPRO current to put the NPRO power back to the level when the MOPA output was higher. But it did not bring back the MOPA power.
So I put back the current after an hour. This caused the sharp power drop on Jan26.
By mistake, I did not fully recover the current at that time and left it like that for a day. This accounts for the long power drop period continued until Jan27.

Shortly after I tweaked the current, the MOPA output power started to fluctuate a lot. This drives the ISS crazy.
To see if this was caused by the NPRO or power amplifier,
we decided to fix the 126MON to monitor the real NPRO power.
We opened the MOPA box and installed a mirror to direct a picked off NPRO beam to the outside of the box through an unused hole.
We set up a lens and a PD outside of the MOPA box to receive this beam. The output from the PD is connected to the 126MON cable.
So 126MON is now serving as the real monitor of the NPRO power. It has not yet been calibrated.

The second attachment shows a short time series of the MOPA power and NPRO power. When the beam is blocked, the 126MON goes to -22.
So the RIN of the NPRO is less than 1%, whereas the MOPA power fluctuates about 5%. There is also no clear correlation between the power fluctuation of the MOPA and the NPRO. So probably the MOPA power fluctuation is not caused by NPRO.

At this moment, all the feedback signals (current shunt, slow and fast actuators) are physically disconnected from MOPA box so that we can see the behavior of MOPA itself.
Attachment 1: Recent10Days.png  195 kB  | Hide | Hide all
Recent10Days.png
Attachment 2: 126_MOPA.png  321 kB  | Hide | Hide all
126_MOPA.png
ELOG V3.1.3-